write an equation for the problems and solve : In a class of 42 students, the number of boys is 12 fewer than twice the number of girls. Find the number of boys and girls.
Answer by Theo(13342)

leb b represent the number of boys.
let g represent the number of girls.

your formulas are:
b = 2g - 12
b + g = 42

replace b with 2g - 12 to get:
2g - 12 + g = 42
simplify to get:
3g - 12 = 42
add 12 to both sides of the equation to get:
3g = 54
divide both sides of the equation by 3 to get:
g = 18
since b = 2g - 12, this means that:
b = 24

you have 24 boys and 18 girls.
the total is 42 students.
the number of boys is equal to 12 fewer than 2 times the girls.
this is because 2 times the girls is 36 and 12 fewer than that is 24.